OpenStudy (anonymous):

What is the slope of a line that passes through the points (8, –3) and (–2, 4)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

I'm not even sure if this is right. M= -7/6?

hartnn (hartnn):

-7 is correct! not 6 \(2- (-8) = 2+8 = ..?\)

OpenStudy (anonymous):

Okay thanks but that isn't one of the options. I have to choose from -10 over 7, -7 over ten, one over ten, and -1 over 6.

hartnn (hartnn):

2+8 =10 -7/10 means -7 over 10 .-.

OpenStudy (anonymous):

Okay, thank you so much. I'm like....really really bad at math!

hartnn (hartnn):

no, you're not! trust me :)

OpenStudy (anonymous):

You are a life saver though.

hartnn (hartnn):

oh, that i am, hehe, thanks :D